OpenStudy (anonymous):

the area rectangle portion of a national park is (48x^2+16x) square km.if the lenght is 8x km long,what is the width,in km? please help me

OpenStudy (anonymous):

It's like the last one we did. Divide EACH term of the AREA by the length.

OpenStudy (anonymous):

\[48x ^{2}/8x \] and \[16x/8x\]

OpenStudy (anonymous):

Then add your two results together (you won't be able to "combine" them because they aren't LIKE TERMS. But you can put them side by side.)

OpenStudy (anonymous):

What'd you get?

OpenStudy (anonymous):

its hard and my first time doing it

OpenStudy (espex):

Maybe stating it a different way will help. You know that: Area = Width * Length \[Area = 48x^2 + 16x\] Length = 8x Width = ? So you have \[48x^2+16x = 8x\space \times \space ?\] Solve for your ? First thing you would do is to divide both sides by 8x \[\frac{48x^2 + 16x}{8x} = \frac{8x \times ?}{8x}\] Leaving you with \[\frac{48x^2 + 16x}{8x} = ?\] Break it into separate fractions \[\frac{48x^2}{8x} + \frac{16x}{8x} = ?\] Solve and your answer will be the width you are looking for.

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

6x + 3